OpenStudy (beccab003):

if a number is -6/0 is it undefined? Also when it is opposite like: 0/-6 the answer is 0. Is that correct?

OpenStudy (solomonzelman):

yes, that is correct.

OpenStudy (solomonzelman):

Whenever you have a zero in a denominator the answer is undefined, whenever you have a zero in numerator the answer is zero.

OpenStudy (beccab003):

Great! thanks!!

OpenStudy (solomonzelman):

No problem... Need help come here more. ALSO.... Also, for a future reference: you can find any section by: 1) clicking "more subjects" ( top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying. you can fan someone by: 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" ) you can mention someone in your question (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @BeccaB003 (they would get a notification that they got mentioned just like you did) you can message someone by: 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. If you have any questions you can message me or anyone else on this website. We will be happy to respond !

OpenStudy (beccab003):

Thanks for that info!! Are you a person who started this site?

OpenStudy (solomonzelman):

NOoo :) Preetha is the person who started it. I am here for 126 days...

OpenStudy (beccab003):

okay. Can you help me with this question?

OpenStudy (beccab003):

Choose the correct equation of a line, in point-slope form, given the point (-1,3) and a slope, m, of 1/2. 1) y+1=1/2(x-3) 2) y-3=1/2(x+1) 3) y+3=1/2(x-1) 4) y-1=1/2(x-3)